Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 12:29 Friday 12 March 2021 (16246) Print this report
Comment to Test of PR3 mirror pitch stability using the Donald Trump method. (16239)

With Hirata-san

See pictures in album PR3 Remedying Work.

After subjecting the suspension to the Trump procedure, today we inspected it visually just to make sure everything was well.

  • We did not find anything wrong.
  • The only change was the rotation of the H2 OSEM flag, which can be seen from before and after pictures. We haven't fixed it yet.
  • We also inspected the wire breakers. See pictures with today's date in the album.
  • We inspected the places where we put the adhesive on the wire. We did not notice any obvious change. Again see pictures with today's date in the album. 

At the end of the session we turned on the oplev laser level and checked that the beam was still in the same place. It was indeed, namely close to 103 mm. See yesterday's entry  notebook entry here.

Search Help